# Emergent classical gravity as stress-tensor deformed field theories

### v1

@@ III Linearized Einstein gravity
Higher-order deformations can be obtained perturbatively, as detailed in Supplemental Material A. On the fixed background $\hat{\gamma}$, the stress tensor deformation is complicated and non-local, as the solution to Eq. (8) lacks a compact analytical form. A more practical approach involves separating the metric saddle point from the field theory action and analyzing the effective field theory on the deformed metric.
In the formulation of effective deformed action (5), substituting Eq. (8) into Eq. (7) yields $\mathcal{L}^{(\lambda)}_{\text{st}}=\frac{\sqrt{g^{*}}}{d-2}(\hat{T}^{*}+\frac{2l^{2}\Lambda}{\lambda})$. The corresponding flow equation follows immediately,
$\partial_{\lambda}S_{\mathrm{EFT}}^{(\lambda)}=-\frac{2l^{2}\Lambda\,\mathrm{Vol}(\mathcal{M})}{(d-2)\lambda^{2}}.$ (12)
The effective stress tensor is obtained by taking the functional derivative of $S_{\text{EFT}}$ with respect to the metric,
$T_{\mathrm{EFT}}^{(\lambda)\mu\nu}\!=\!\hat{T}^{\mu\nu}\!-\!\frac{1}{d\!-\!2}\left(\!\hat{T}\!+\!\frac{2l^{2}\Lambda}{\lambda}\!\right)g^{\mu\nu}\!-\!\frac{2}{d\!-\!2}\frac{\partial\hat{T}}{\partial g_{\mu\nu}}.$ (13)
which depends on the explicit form of $\hat{T}$. For a massless free scalar seed, $\hat{\mathcal{L}}=\frac{\sqrt{g}}{2}\nabla^{\mu}\phi\nabla_{\mu}\phi$, one has $g_{\mu\nu}\,\frac{\partial\hat{T}}{\partial g_{\mu\nu}}=\hat{T}$. By taking the trace of Eq. (13), $\hat{T}$ can be rewritten in terms of $T_{\text{EFT}}$. Combining this with Eq. (6) yields the total $\lambda$-derivative of the effective field action,
$\frac{\text{d}}{\text{d}\lambda}\!S_{\mathrm{EFT}}^{(\lambda)}[g^{*}\!,\psi]\!=\!\frac{1}{4\lambda}\!\int\!d^{d}x\sqrt{\!g^{*}}\!\left(\!T_{\mathrm{EFT}}^{(\lambda)*}\!+\!\frac{2(d\!-\!4)l^{2}\Lambda}{(d\!-\!2)\lambda}\!\right).$ (14)
On the deformed metric, the flow equation of the effective field theory action is both compact and local in form. This approach admits a systematic extension to a broad class of gravitational theories.
IV Palatini $f(\mathcal{R})$ gravity
In this section, the Palatini $f(\mathcal{R})$ gravity is investigated, corresponding to the field theory deformation induced by the trace of the stress tensor. In the Palatini formalism, the variations of the metric and the connection are independent. The Riemann curvature tensor $\mathcal{R}^{\mu}{}_{\nu\rho\sigma}$ and the Ricci curvature tensor $\mathcal{R}_{\mu\nu}$ are constructed with the independent connection. The gravitational Lagrangian is formally written as $\mathcal{L}^{(\lambda)}_{\text{grav}}=\frac{\sqrt{g}}{2\lambda}f(\mathcal{R})$, where the Ricci scalar $\mathcal{R}=g^{\mu\nu}\mathcal{R_{\mu\nu}}$. Varying with respect to $g^{\mu\nu}$ and $\Gamma^{\sigma}{}_{\mu\nu}$ yields the equations of motion:
$\begin{split}f^{\prime}(\mathcal{R})\mathcal{R}_{(\mu\nu)}-\frac{1}{2}f(\mathcal{R})g_{\mu\nu}&=-\lambda\hat{T}_{\mu\nu},\\ \bar{\nabla}_{\sigma}(\sqrt{g}f^{\prime}(\mathcal{R})g^{\mu\nu})&=0.\end{split}$ (15)
Here $\bar{\nabla}$ denotes the covariant derivative with respect to $\Gamma^{\sigma}{}_{\mu\nu}$. For a certain $f$, the trace of the first EOM in (15) provides an algebraic equation between $\mathcal{R}$ and $\hat{T}$. The second EOM in (15) can be solved by introducing the following metric [20], $\tilde{g}_{\mu\nu}=(f^{\prime}(\mathcal{R}))^{\frac{2}{d-2}}g_{\mu\nu}$, and the solution $\Gamma^{\sigma}{}_{\mu\nu}$ is the Levi-Civita connection of $\tilde{g}_{\mu\nu}$. Subsequently, (15) reduces to Einstein’s equation featuring a modified stress tensor [20], which can be solved by applying the perturbation method presented in the previous section. A detailed discussion is provided in Supplementary Material B.
IV.1 Four-dimensional quadratic Palatini gravity
As an illustrative example, the scenario with four-dimensional spacetime and a quadratic form of the function $f$ is considered,
$S^{(\lambda)}_{\text{grav}}[g,\Gamma]$ $=\frac{l^{2}}{2\lambda}\int d^{4}x\sqrt{g}\Big{(}-2\Lambda+\mathcal{R}+\alpha l^{2}\mathcal{R}^{2}\Big{)}.$ (16)
With $\Lambda=0$ and $\alpha=\kappa/(2l^{2})$, the action reduces to the Starobinsky model [27]. Taking the trace of Eq. (15) and substituting it into the action yields the effective deformed action:
$S^{(\lambda)}_{\text{EFT}}[g^{*},\psi]$ $=\hat{S}[g^{*},\psi]+\int d^{4}x\sqrt{g^{*}}\Big{[}\frac{\alpha\lambda}{2}\Big{(}\hat{T}^{*}\Big{)}^{2}$
$\quad+\frac{1+8\alpha l^{2}\Lambda}{2}\Big{(}\hat{T}^{*}+\frac{2\Lambda}{\lambda l^{-2}}\Big{)}\Big{]}.$ (17)
The effective stress tensor $T_{\mathrm{EFT}}$ depends on the explicit form of $\hat{T}$. For concreteness, we take a massless scalar seed with $\hat{T}_{\mu\nu}=\nabla_{\mu}\phi\,\nabla_{\nu}\phi-\tfrac{1}{2}g_{\mu\nu}(\nabla\phi)^{2}$. Varying Eq. (17) with respect to $g_{\mu\nu}$ and inverting expresses $\hat{T}$ in terms of $T_{\mathrm{EFT}}$, from which the flow equation follows:
$\partial_{\lambda}S^{(\lambda)}_{\text{EFT}}$ $\!=\!\int\!d^{4}x\sqrt{g}\Big{[}-\frac{l^{2}\Lambda(1\!+\!8\alpha l^{2}\Lambda)}{\lambda^{2}}$
$\quad+\!\frac{1}{2\alpha}\Big{(}\frac{\!1\!+\!12\alpha l^{2}\Lambda\!-\!\sqrt{\!(1\!+\!4\alpha l^{2}\Lambda)^{2}\!-\!4\alpha\lambda T^{(\lambda)}_{\text{EFT}}}}{4\lambda}\Big{)}^{2}\Big{]}.$ (18)
The total derivative of the effective deformed action is given by
$\frac{\text{d}}{\text{d}\lambda}S^{(\lambda)}_{\text{EFT}}[g^{*},\psi]$ $\!=\!\frac{1}{8\lambda}\!\int\!d^{4}x\sqrt{g^{*}\!}\Big{[}T^{(\lambda)}_{\text{EFT}}\!+\!\frac{1\!+\!8\alpha l^{2}\Lambda(1\!-\!2\alpha l^{2}\Lambda)}{2\alpha\lambda}$
$\quad-\!\frac{(1\!+\!4\alpha l^{2}\Lambda)\sqrt{\!(1\!+\!4\alpha l^{2}\Lambda)^{2}\!-\!4\alpha\lambda T^{(\lambda)}_{\text{EFT}}}}{2\alpha\lambda}\Big{]}.$ (19)
V Emergent non-minimal gravity
@@
We proceed to examine a more general action, in which the metric, curvature, and matter fields exhibit non-minimal coupling within the Lagrangian. Formally, this action  [^1] Here we assume that the action does not include the Riemann curvature tensor $\mathcal{R^{\mu}_{\nu\rho\sigma}}$. In principle, $\mathcal{A}^{(\mu)}$ should be a local function of all independent invariants composed of $g^{\mu\nu}$, $\mathcal{R}_{\mu\nu}$, and $X^{(i)}_{\mu\nu}$. can be written as
$S^{(\lambda)}_{\text{grav}}[g,\psi]$ $=\int d^{d}x\sqrt{g}\mathcal{A}^{(\lambda)}(g^{\mu\nu},\mathcal{R}_{\mu\nu},X^{(i)}_{\mu\nu},\phi^{(i)}),$ (20)
@@
where $\mathcal{R}_{\mu\nu}$ is the Ricci curvature tensor in the Palatini formalism, and the tensor $X^{(i)}_{\mu\nu}=\nabla_{\mu}\phi^{(i)}\nabla_{\nu}\phi^{(i)}$. Variations of (20) with respect to the metric and the independent connection yield
$\begin{split}\frac{\partial\mathcal{A}^{(\lambda)}}{\partial g^{\mu\nu}}-\frac{1}{2}\mathcal{A}^{(\lambda)}g_{\mu\nu}&=0,\\ \bar{\nabla}_{\sigma}\Big{(}\sqrt{g}\frac{\partial\mathcal{A}^{(\lambda)}}{\partial\mathcal{R}_{\mu\nu}}\Big{)}&=0.\end{split}$ (21)
Substituting the first equation into the action and eliminating $\mathcal{R}_{\mu\nu}$ yields
$S_{\text{grav}}^{(\lambda)}[g^{*},\psi]$ $=\int d^{d}x\sqrt{g}\mathcal{A}^{(\lambda)}(g^{\mu\nu},\mathcal{R}_{\mu\nu},X^{(i)}_{\mu\nu},\phi^{(i)})\Big{|}_{g=g^{*}}$
$=\int d^{d}x\sqrt{g^{*}}\mathcal{B}^{(\lambda)}(g^{*\mu\nu},X^{(i)}_{\mu\nu},\phi^{(i)}).$ (22)
Following the prescription in the previous section, the effective deformed action is defined by $S^{(\lambda)}_{\text{EFT}}[g,\psi]=\int d^{d}x\sqrt{g}\mathcal{B}^{(\lambda)}(g^{\mu\nu},X^{(i)}_{\mu\nu},\phi^{(i)})$. The flow equation is
$\partial_{\lambda}S^{(\lambda)}_{\text{EFT}}$ $=\int d^{d}x\sqrt{g}\partial_{\lambda}\mathcal{B}^{(\lambda)}.$ (23)
Assuming that the theory contains only a single tensor field defined by $X_{\mu\nu}=\sum_{i,j}G_{ij}(\phi)\nabla_{\mu}\phi^{(i)}\nabla_{\nu}\phi^{(j)}$. There exist $d$ independent invariants constructed from $X_{\mu\nu}$, which can be expressed as $X_{n}=X^{\mu_{n}}_{\mu_{1}}X^{\mu_{1}}_{\mu_{2}}\cdots X^{\mu_{n-1}}_{\mu_{n}}=\text{tr}(X^{n})$, for n=1,2,,dn=1,2,...,d. The Lagrangian $\mathcal{B}^{(\lambda)}$ is a local function of $X_{n}$ and $\phi$. Thus, the effective stress tensor derived from (22) takes the form
$(T^{(\lambda)}_{\text{EFT}})^{\mu}_{\nu}$ $=2\sum_{n=1}^{d}n(X^{n})^{\mu}_{\nu}\partial_{X_{n}}\mathcal{B}^{(\lambda)}-\mathcal{B}^{(\lambda)}\delta^{\mu}_{\nu}.$ (24)
Suppose that $X^{\mu}_{\nu}$ is diagonalizable, expressed as $X=U\text{diag}(\chi_{1},\chi_{2},...,\chi_{d})U^{-1}$. According to (24), the effective stress tensor can be diagonalized by the same matrix $U$, and its eigenvalues $\{{\tau_{j}}\}$ are given by
$\tau_{j}^{(\lambda)}$ $=2\chi_{j}\partial_{\chi_{j}}\mathcal{B}^{(\lambda)}-\mathcal{B}^{(\lambda)},\ \ \text{for}\ \ j=1,2,...,d.$ (25)
For a given $\mathcal{B}^{(\lambda)}$, Eq. (25) can be inverted to express $\chi_{j}$ in terms of $\tau_{j}$; inserting this into Eq. (23) yields a flow equation written entirely in the effective stress tensor.
V.1 Example: Generalized Nambu-Goto action
$\quad S^{(\lambda)}_{\text{EFT}}\!=\!\int\!d^{d}x\sqrt{g}\Big{[}\frac{1\!-\!2\lambda V\!-\!\sqrt{1\!-\!2\lambda(1\!-\!\lambda V)\nabla^{\mu}\phi\nabla_{\mu}\phi}}{\lambda(1\!-\!\lambda V)}\Big{]},$ (26)
$\partial_{\lambda}S^{(\lambda)}_{\text{EFT}}$ $=\!\int\!d^{d}x\sqrt{g}\Big{[}\frac{1}{2d}\text{tr}(T^{2}_{\text{EFT}})\!-\!\frac{1}{d^{2}}(\text{tr}T_{\text{EFT}})^{2}$
$\quad-\frac{d\!-\!2}{2d^{3/2}\sqrt{d\!-\!1}}\text{tr}T_{\text{EFT}}\sqrt{\text{tr}(T^{2}_{\text{EFT}})\!-\!\frac{1}{d}(\text{tr}T_{\text{EFT}})^{2}}\Big{]}.$ (27)
From the perspective of gravity, we propose the following action in the Palatini formalism (for $d\geq 3$),
$\mathcal{A}^{(\lambda)}$ $\!=\!\frac{2(d\!-\!1)(1\!-\!\lambda V)\nabla^{\mu}\phi\nabla_{\mu}\phi}{f(V)}\!-\!\frac{(d\!-\!2)(1\!-\!2\lambda V)}{2\lambda(1\!-\!\lambda V)}$
$\quad+\!\frac{f(V)}{4\lambda(d\!-\!1)}\Big{(}\!\frac{(d\!-\!1)(d\!-\!2)}{(1\!-\!\lambda V)^{2}}\!+\!\frac{l^{2}\mathcal{R}}{(d\!-\!2)^{2}}\!-\!\frac{2l\sqrt{\mathcal{R}}}{1\!-\!\lambda V}\!\Big{)},$ (28)
where the function $f(V)=d(1-\lambda V)(1-2\lambda V)-\sqrt{(1-\lambda V)^{2}(4-4d+d^{2}(1-2\lambda V)^{2})}$. By utilizing the first EOM in (21), the curvature terms present in (28) are eliminated, thereby recovering the generalized Nambu-Goto action (26). In particular, when $V(\phi)=0$, the gravitational action (28) reduces to the minimal coupling between a free scalar field and the Palatini $f(\mathcal{R})$-gravity,
The gravitational saddle satisfies the Einstein equation with a modified stress tensor, solvable via the perturbation method in Supplemental Material A. More generally, by extending the form of the effective deformed action (26), the gravitational action $\mathcal{A}^{(\lambda)}$ reduces to Einstein or quadratic Palatini action under specific conditions, as detailed in Supplemental Material C.
V.2 Example: $T\bar{T}$-like deformation in $d$ dimensions
@@
where $\{{\beta_{j}}\}$ is the deformation parameters of the root-$T\bar{T}$-like operator as discussed in [18], $\{{p_{j}}\}$ are the numbers that characterizing the deformation, $\Sigma=\sum_{j=1}^{d}1/p_{j}$, and $\Delta=(2\Sigma+d-4)d/2$. $\mathcal{B}_{0}$ satisfies the differential equation $2\chi_{j}\partial_{\chi_{j}}\mathcal{B}_{0}-\mathcal{B}_{0}=0$. The solution takes the form
$\mathcal{B}_{0}=C(\phi)\prod_{j=1}^{d}\chi_{j}^{\frac{1}{2}}=C(\phi)\frac{\sqrt{\det(X_{\mu\nu})}}{\sqrt{\det(g_{\mu\nu})}},$ (31)
which is analogous to the action of the Nambu–Goto string [30]. The eigenvalues of the effective stress-energy tensor $\{\tau^{(\lambda)}_{j}\}$ are obtained by plugging (30) into (25). These eigenvalues can then be substituted into (23) to derive the corresponding flow equation,
$\partial_{\lambda}S_{\text{EFT}}^{(\lambda)}$ $\!=\!(1-\Sigma)l^{\frac{\Delta}{1\!-\!\Sigma}}b^{\frac{1}{2(1-\Sigma)}}\!\int\!d^{d}x\sqrt{g}\Big{(}\!\prod_{j=1}^{d}(\tau_{j}^{(\lambda)})^{\frac{1}{p_{j}}}\!\Big{)}^{\frac{1}{\Sigma-1}},$ (32)
where $b=\prod_{j=1}^{d}\beta_{j}$. In particular, by setting $p_{1}=p_{2}=\dots=p_{d}=p$, the deformation operator reduces to the determinant of the effective stress tensor, $\mathcal{O}_{\text{st}}=(\det[(T^{(\lambda)}_{\text{EFT}})^{\mu}_{\nu}])^{\frac{1}{d-p}}$ [16, 14].
Next, a class of gravitational actions is constructed that reproduces the effective deformed action given in (30). For simplicity, let the Lagrangian $\mathcal{A}^{(\lambda)}$ to be a local function of the invariants $X_{n}$ and those built from the Ricci tensor, $\mathcal{R}_{n}=\text{tr}(\mathcal{R}^{n})$. Suppose that $\mathcal{R}^{\mu}_{\nu}$ is diagonalizable and can be expressed as $\mathcal{R}=\tilde{U}\,\text{diag}(\tilde{r}_{1},\tilde{r}_{2},...,\tilde{r}_{d})\tilde{U}^{-1}$. Under these assumptions, the first EOM in (21) reduces to
$U^{-1}\tilde{U}\text{diag}(2\tilde{r}_{1}\partial_{\tilde{r}_{1}}\mathcal{A}^{(\lambda)},...,2\tilde{r}_{d}\partial_{\tilde{r}_{d}}\mathcal{A}^{(\lambda)})(U^{-1}\tilde{U})^{-1}$
$=\text{diag}(\mathcal{A}^{(\lambda)}\!-\!2\chi_{1}\partial_{\chi_{1}}\mathcal{A}^{(\lambda)},...,\mathcal{A}^{(\lambda)}\!-\!2\chi_{d}\partial_{\chi_{d}}\mathcal{A}^{(\lambda)}).$ (33)
The two matrices have identical diagonal entries up to a permutation. Relabel the eigenvalues of $\mathcal{R}^{\mu}_{\nu}$ as $\{r_{j}\}$, which satisfy $2r_{j}\partial_{r_{j}}\mathcal{A}^{(\lambda)}=\mathcal{A}^{(\lambda)}-2\chi_{j}\partial_{\chi_{j}}\mathcal{A}^{(\lambda)}$. The explicit form of $\mathcal{A}^{(\lambda)}$ is introduced as follows:

### v2

@@ III Linearized Einstein gravity as stress tensor deformed theory
The leading deformation organizes as a spin-2 channel exchange controlled by $\hat{\square}^{-1}$, providing a concrete check against the linearized Einstein response. The resulting deformation (10) is intrinsically nonlocal on a reference background $\hat{\gamma}_{\mu\nu}$ and reproduces the linearized Einstein response for both scalar and Maxwell theories (SM A). Higher-order corrections can be computed systematically in perturbation theory (SM A); nonlocality is inevitable at every order in the deformation  [^5] Since the deformation generates the intrinsically nonlocal kernel (10), the locality assumptions entering the Weinberg–Witten theorem do not apply, and the saddle-point correspondence to Einstein dynamics is therefore not in conflict with that theorem. Once the Green’s function is specified, the operator $1/\Box$ is fixed up to local counterterms; in Minkowski signature, adopting the retarded prescription enforces causality. . Moreover, for Gauss-Bonnet gravity, the leading-order deformation retains the same functional form as (III), with the reference metric $\hat{\gamma}_{\mu\nu}$ replaced by the corresponding Gauss–Bonnet background. The same saddle-point analysis continues to apply; details are presented in SM A.
In the formulation of effective deformed action (5), substituting (8) into (7) yields $\mathcal{L}^{(\lambda)}_{\text{st}}=\frac{\sqrt{g^{*}}}{d-2}\hat{T}^{*}$. The effective stress-tensor is obtained by taking the functional derivative of $S_{\text{EDA}}$ with respect to the metric, $T_{\mathrm{EDA}}^{(\lambda)\mu\nu}=\hat{T}^{\mu\nu}-\frac{1}{d-2}\hat{T}g^{\mu\nu}-\frac{2}{d-2}\frac{\partial\hat{T}}{\partial g_{\mu\nu}}$. For a massless free scalar seed, $\hat{\mathcal{L}}=\frac{\sqrt{g}}{2}\nabla^{\mu}\phi\nabla_{\mu}\phi$, one has $g_{\mu\nu}\,\frac{\partial\hat{T}}{\partial g_{\mu\nu}}=\hat{T}$. The trace of the seed theory stress-tensor $\hat{T}$ can be rewritten in terms of $T_{\text{EDA}}$. Combining this with (6) yields the total $\lambda$-derivative of the effective field action, $\frac{\text{d}}{\text{d}\lambda}S_{\mathrm{EDA}}^{(\lambda)}[g^{*},\psi]=\frac{1}{4\lambda}\int d^{d}x\sqrt{\!g^{*}}T_{\mathrm{EDA}}^{(\lambda)*}$. In four-dimensional Maxwell theory the traceless stress tensor preserves conformal invariance, and the linearized Einstein–Hilbert deformation produces the nonlocal action (III) whose surviving transverse–traceless spin-2 exchange reproduces the standard graviton-mediated backreaction, demonstrating that the metric saddle responds nonlocally to the seed stress tensor and thereby links Einstein dynamics to stress–tensor flows extendable to broader theories.
IV Non-minimal Palatini gravity v.s. $T\bar{T}$-like flow
Having established the framework and its realizations in Einstein gravity, we next address the inverse problem: at the classical (saddle) level, we reconstruct non-minimal, metric-affine gravitational actions directly from specified stress-tensor deformations of a seed QFT and verify the correspondence on representative cases via their flow equations. Our discussion is based on the Palatini formalism, in which the Riemann tensor $\mathcal{R}^{\mu}{}_{\nu\rho\sigma}$ and the Ricci tensor $\mathcal{R}_{\mu\nu}$ are constructed with the independent connection. Starting from a specific effective deformed action defined as $S^{(\lambda)}_{\text{EDA}}[g,\psi]=\int d^{d}x\sqrt{g}\mathcal{B}^{(\lambda)}(g^{\mu\nu},X_{\mu\nu},\psi)$, where the tensor $X_{\mu\nu}=\sum_{i,j}\mathcal{G}_{ij}(\psi)\nabla_{\mu}\phi^{(i)}\nabla_{\nu}\phi^{(j)}$ and $\mathcal{G}_{ij}$ is the target space metric. There are $d$ independent invariants constructed from $X_{\mu\nu}$, expressed as $X_{n}=X^{\mu_{n}}_{\mu_{1}}X^{\mu_{1}}_{\mu_{2}}\cdots X^{\mu_{n-1}}_{\mu_{n}}=\text{tr}(X^{n})$ for n=1,2,,dn=1,2,...,d. The Lagrangian $\mathcal{B}^{(\lambda)}$ is a local function of $X_{n}$ and $\phi^{(i)}$, and the effective stress-tensor is given by $(T^{(\lambda)}_{\text{EDA}})^{\mu}_{\nu}=2\sum_{n=1}^{d}n(X^{n})^{\mu}_{\nu}\partial_{X_{n}}\mathcal{B}^{(\lambda)}-\mathcal{B}^{(\lambda)}\delta^{\mu}_{\nu}$.
@@
To derive the gravitational theory corresponding to $S^{(\lambda)}_{\text{EDA}}$, we postulate a candidate gravitational action, which is then verified through its equations of motion to check if it reproduces $\mathcal{B}^{(\lambda)}$. Here, we consider a general class of actions with non-minimal couplings between the metric, curvature, and matter fields. The gravitational action  [^6] Here we assume that the action does not include the Riemann curvature tensor $\mathcal{R^{\mu}_{\nu\rho\sigma}}$. In principle, $\mathcal{A}^{(\mu)}$ should be a local function of all independent invariants composed of $g^{\mu\nu}$, $\mathcal{R}_{\mu\nu}$, and $X^{(i)}_{\mu\nu}$. can be written as:
$S^{(\lambda)}_{\text{grav}}[g,\psi]$ $=\int d^{d}x\sqrt{g}\mathcal{A}^{(\lambda)}(g^{\mu\nu},\mathcal{R}_{\mu\nu},X_{\mu\nu},\psi),$ (11)
@@
where $\mathcal{R}_{\mu\nu}$ is the Ricci tensor in the Palatini formalism. Variations of (11) with respect to the metric and the independent connection yield
$\frac{\partial\mathcal{A}^{(\lambda)}}{\partial g^{\mu\nu}}-\frac{1}{2}\mathcal{A}^{(\lambda)}g_{\mu\nu}$ $=0,$ (12a)
$\bar{\nabla}_{\sigma}\Big(\sqrt{g}\frac{\partial\mathcal{A}^{(\lambda)}}{\partial\mathcal{R}_{\mu\nu}}\Big)$ $=0.$ (12b)
A suitable gravitational Lagrangian $\mathcal{A}^{(\lambda)}$ yields the deformed field theory Lagrangian $\mathcal{B}^{(\lambda)}$ when evaluated at the metric saddle,
$S_{\text{grav}}^{(\lambda)}[g^{*},\psi]$ $=\int d^{d}x\sqrt{g^{*}}\mathcal{B}^{(\lambda)}(g^{*\mu\nu},X_{\mu\nu},\psi).$ (13)
The flow equation of the deformed field theory is given by $\partial_{\lambda}S^{(\lambda)}_{\text{EDA}}[g^{*},\psi]=\int d^{d}x\sqrt{g^{*}}\partial_{\lambda}\mathcal{B}^{(\lambda)}(g^{*\mu\nu},X_{\mu\nu},\psi)$. Different choices of $\mathcal{A}^{(\lambda)}$ lead to different metric saddles $g^{*}$ and hence distinct deformed actions on the reference background. Conversely, given the deformed field theory action on a reference background, the corresponding gravitational action can be reconstructed order by order in metric perturbations. As a realization, we show in SM C that the Generalized Nambu-Goto action arises from a specific Palatini Lagrangian with a non-minimal coupling.
IV.1 Example: $T\bar{T}$-like deformation in $d$ dimensions
@@
where $\{{\beta_{j}}\}$ is the deformation parameters of the root-$T\bar{T}$-like operator as discussed in [6], $\{{p_{j}}\}$ are the numbers that characterizing the deformation, $\Sigma=\sum_{j=1}^{d}1/p_{j}$, and $\Delta=(2\Sigma+d-4)d/2$. $\mathcal{B}_{0}$ satisfies the differential equation $2\chi_{j}\partial_{\chi_{j}}\mathcal{B}_{0}-\mathcal{B}_{0}=0$ and the solution is given by  [^7] The solution takes the form $\mathcal{B}_{0}=C(\phi)\prod_{j=1}^{d}\chi_{j}^{\frac{1}{2}}=C(\phi)\frac{\sqrt{\det(X_{\mu\nu})}}{\sqrt{\det(g_{\mu\nu})}},$ (15) which is analogous to the action of the Nambu-Goto string [31] up to a regular function $C(\phi)$.. The effective stress-tensor can be diagonalized by the same matrix $U$, and its eigenvalues $\{{\tau_{j}}\}$ are given by $\tau_{j}^{(\lambda)}=2\chi_{j}\partial_{\chi_{j}}\mathcal{B}^{(\lambda)}-\mathcal{B}^{(\lambda)}$. These eigenvalues can then be substituted into (14) to derive the flow equation
$\partial_{\lambda}S_{\text{EDA}}^{(\lambda)}$ $\!=\!\frac{(1-\Sigma)}{(l^{\Delta}b^{\frac{1}{2}})^{\frac{1}{\Sigma-1}}}\!\int\!d^{d}x\sqrt{g}\Big(\!\prod_{j=1}^{d}(\tau_{j}^{(\lambda)})^{\frac{1}{p_{j}}}\!\Big)^{\frac{1}{\Sigma-1}},$ (16)
where $b=\prod_{j=1}^{d}\beta_{j}$. In particular, setting $p_{1}=p_{2}=\dots=p_{d}=p$ reduces the deformation operator to $\mathcal{O}_{\text{st}}=(\det[(T^{(\lambda)}_{\text{EDA}})^{\mu}_{\nu}])^{\frac{1}{d-p}}$ [9, 10].
Next, a class of plausible gravitational actions is postulated that could reproduce the effective deformed action given in (14). The explicit form of the gravitational Lagrangian $\mathcal{A}^{(\lambda)}$ is introduced as follows: